# Operational meaning of the fidelity and the Fisher-Bures path length

### v1

@@ Abstract
We show that the minimum entropy production in near-reversible quantum state transport along a path is simple function of the path length measured according to the Fisher–Bures metrics. Consequently, the minimum entropy production along a geodesics is the simple function of fidelity. Hence the sharp values of path lengths, also called statistical lengths, and of fidelities obtain operational meaning to quantify the residual irreversibility in near-reversible state transport. Compared to fidelity’s operative interpretations based on Uhlmann’s theorem, our proposal does not assume extended operations on the purified states. As to the classical limit, the Bhattacharryya fidelity obtains a sharp operational meaning after eighty years.
@@ I Introduction
Fidelities are related to distances, also called Fisher–Bures statistical distances, on the underlying Riemannian geometries. The infinitesimal distance $d\ell$ defines the Fisher metrics Fisher (1925) between the probability distributions $p$ and $p+dp$ and the Bures metrics Bures (1969) between the quantum states $d\hat{\rho}$ and $\hat{\rho}+d\hat{\rho}$:
@@
$\ell(p.q)_{Fisher}$ $=$ $2\arccos F(p,q)~{}\in[0,\pi],$ (5)
$\ell(\hat{\rho}.\hat{\sigma})_{Bures}$ $=$ $2\sqrt{1-F^{2}(\hat{\rho},\hat{\sigma})}~{}\in[0,2].$ (6)
@@
The present work proposes the operational meaning of path lengths in the Fisher–Bures metrics, which implies the meaning of the geodesics and the fidelities. Differently from ref. Dodd and Nielsen (2002), we are not relying on Uhlmann’s theorem and the ‘environmental’ system. The concept is that in near-reversible transport along a path $\gamma$ between two predefined states $\hat{\rho},\hat{\sigma}$ the minimum entropy production is quantified by the length $\ell_{\gamma}(\hat{\rho},\hat{\sigma})$.
@@ II State transport by equilibrating reservoirs
The entropy production can be smaller and can even go to zero if we apply sequential equilibration with intermediate reservoirs along a path, such that the stepsizes go to zero. The following feature of relative entropy will be important:
@@
where $d\ell$ is defined by the Bures metrics (4). Consider a smooth (not necessarily geodesic) path $\gamma$ from $\hat{\rho}$ to $\hat{\sigma}$. We use $N$ reservoirs to transport $\hat{\rho}$ into $\hat{\sigma}$ in $N$ steps. Consider a monotone sequence of intermediate states $\hat{\rho}_{i}$ along the path, such that $\hat{\rho}=\hat{\rho}_{0}$ and $\hat{\rho}_{N}=\hat{\sigma}$. Perform the transports $\hat{\rho}\Rightarrow\hat{\rho}_{1}\Rightarrow\hat{\rho}_{2}\Rightarrow\dots\Rightarrow\hat{\rho}_{N-1}\Rightarrow\hat{\sigma}$. The total entropy production is the sum of the $N$ yields:
@@
The fastest approach to reversible transport is achieved on geodesics. Inserting the geodesic lengths (5-6) in (15) and its classical counterpart, the ultimate lower bounds on entropy production read
$\Delta S=\frac{2}{N}\left\{\begin{array}[]{ll}1-F^{2}&\mbox{(quantum)}\\ (\arccos F)^{2}&\mbox{(classical)}\end{array}\right.$ (16)
These express the desired sharp operational meaning of both the classical and quantum fidelities (1-2). The bounds are attainable in the limit $N\rightarrow\infty$ by operations on the system plus the reservoirs.
We find the simple meaning of the Bures length if we introduce the number $\nu$ of equilibrations per unit length, i.e., the density $\nu=N/\ell_{\gamma}$. Then the entropy production becomes the following linear function of the path length:
@@
This suggests the following operational meaning of the Bures length. Consider the near-reversible state transport via sequential equilibrations along a smooth path. If $\nu\gg 1$ is the density of equilibrations per unit path length then the entropy production is $(1/2\nu)$ per unit length.
@@ III Remarks and summary
Paths in Fisher–Bures geometries are perhaps the simplest non-local objects and, as such, they can be interpreted as state transport processes. It has therefore seemed logical to look here for the meaning of the geometries. One should mention the standard Kantorovich–Wasserstein theory Kantorovich (1960); Vaserstein (1969) where the transport from distribution $p$ to $q$ happens via direct relocation of populations from the $p_{\alpha}$’s into the $q_{\beta}$’s. The transport of quantum state $\hat{\rho}$ into $\hat{\sigma}$, too, can work similarly after a single unitary rotation of $\hat{\rho}$ that makes it commute with $\hat{\sigma}$. This may seem much simpler than sequential equilibration. But the simple method assumes the detailed control of the states in question. If it is not possible because, e.g., the states are complicated many-body states then the method of equilibration becomes more valuable.
@@
Thermodynamic systems are the typical examples. The Weinhold–Ruppeiner geometry Weinhold (1975); Ruppeiner (1979) defines the thermodynamic distances between thermodynamic states and the thermodynamic path lengths in general. It has long been known that the thermodynamic lengths coincide with statistical lengths Diósi et al. (1984) and they are rooted in the the Fisher geometry of Gibbs states Janyszek and Mrugała (1989). In thermodynamics, the expression $(\Delta\ell)^{2}/V$ quantifies how the statistical fluctuations of the thermodynamic parameters tend to zero in the thermodynamic limit of infinite volume $V\rightarrow\infty$. This does not interpret the global thermodynamic lengths $\ell$ but the local metrics. Studies of thermodynamic state transport Salamon and Nulton (1998); Salamon et al. (2023); Diósi (2002); Diosi et al. (1996) targeted a relationship between the thermodynamic path length and the thermodynamic entropy production. Salamon and Nulton Salamon and Nulton (1998) recognized that in discrete sequential equilibration the equilibration rates cancel from the expression of entropy production and they found the thermodynamic predecessor of eq. (15). Ref. Diósi (2002) was the first attempt at constructing an underlying classical microscopic mechanism of a specific irreversible state transport. The present work constructed the abstract microscopic generalization valid in the Fisher–Bures geometries of arbitrary quantum states including the classical special case of probability distributions. For this construction, the analytically tractable microscopic reservoir model and the related theorem, both proposed in ref. Diósi et al. (2006), have been instrumental.
@@
The sharp, not just qualitative, operational (physical) meaning of classical and quantum fidelities has been demonstrated in the near-reversible transport of states. On the one hand, classical fidelity and the global Fisher statistical distance got an operational interpretation for the first time. On the other hand, quantum fidelity and Bures distance got an operational interpretation that works without operations on Uhlmann’s extension beyond the system of interest. The concrete choice of the reservoir and the equilibration protocol is not likely to be critical and alternative microscopic models should lead to the same relationships between entropy production and the Fisher–Bures geometry. Future investigations may replace our stepwise transport with some continuous protocol.
@@ Acknowledgements.
This research was funded by the National Research, Development and Innovation Office (Hungary) Frontline Research Excellence Program (Grant No. KKP133827).

### v2

# Operational meaning of the classical fidelity and the path length in Fisher-Kubo-Mori-Bogoliubov geometry

@@ Abstract
We show that the minimum entropy production in near-reversible quantum state transport along a path is simple function of the path length measured according to the Fisher–KMB metrics. Hence the sharp values of path lengths, also called statistical lengths, obtain operational meaning to quantify the residual irreversibility in near-reversible state transport. In the classical limit, the Bhattacharryya fidelity obtains a sharp operational meaning after eighty years.
@@ I Introduction
Fidelities are related to distances, also called statistical distances, on the underlying Riemannian geometries. The infinitesimal distance $d\ell$ defines the Fisher metrics Fisher (1925) between the probability distributions $p$ and $p+dp$ and the Bures metrics Bures (1969) between the quantum states $d\hat{\rho}$ and $\hat{\rho}+d\hat{\rho}$:
@@
$\ell(p,q)_{Fisher}$ $=$ $2\arccos F(p,q)~{}\in[0,\pi],$ (5)
$\ell(\hat{\rho},\hat{\sigma})_{Bures}$ $=$ $2\sqrt{1-F^{2}(\hat{\rho},\hat{\sigma})}~{}\in[0,2].$ (6)
@@
While in metrology and informatics a natural quantum generalization of the Fisher metrics is the Bures one, in irreversible processes it is the Kubo-Mori-Bogoliubov (KMB) metrics (cf., e.g., ref. Hayashi (2002)):
$(d\ell)_{KMB}^{2}=\mathsf{tr}\left(d\hat{\rho}d\!\log\hat{\rho}\right)=\mathsf{tr}\left(d\hat{\rho}\frac{\log\hat{\rho}_{L}-\log\hat{\rho}_{R}}{\hat{\rho}_{L}-\hat{\rho}_{R}}d\hat{\rho}\right).$ (7)
Unlike the expression of geodesic distance in terms of fidelity in Fisher–Bures metrics, no such closed expression is known in the Fisher–KMB geometry. Both geometries have been missing their global interpretations but the close relationship of the KMB metrics with von Neumann entropies will help us. The present work proposes the operational meaning of path lengths in the Fisher–KMB metrics, which implies the meaning of the geodesics. In the special case we obtain the operational meaning of the classical Bhattacharyya fidelity as well. The concept is that in near-reversible transport along a path $\gamma$ between two predefined states $\hat{\rho},\hat{\sigma}$ the minimum entropy production is quantified by the length $\ell_{\gamma}(\hat{\rho},\hat{\sigma})$.
@@ II State transport by equilibrating reservoirs
The entropy production can be smaller and can even go to zero if we apply sequential equilibration with intermediate reservoirs along a path, such that the stepsizes go to zero. The following feature of relative entropy will be important Petz and Toth (1993):
@@
where $d\ell$ is defined by the Fisher metrics (3) in the classical limit and in the general quantum case it corresponds to the Fisher–KMB (7) and not to the Fisher–Bures (4) metrics. Consider a smooth (not necessarily geodesic) path $\gamma$ from $\hat{\rho}$ to $\hat{\sigma}$. We use $N$ reservoirs to transport $\hat{\rho}$ into $\hat{\sigma}$ in $N$ steps. Consider a monotone sequence of intermediate states $\hat{\rho}_{i}$ along the path, such that $\hat{\rho}=\hat{\rho}_{0}$ and $\hat{\rho}_{N}=\hat{\sigma}$. Perform the transports $\hat{\rho}\Rightarrow\hat{\rho}_{1}\Rightarrow\hat{\rho}_{2}\Rightarrow\dots\Rightarrow\hat{\rho}_{N-1}\Rightarrow\hat{\sigma}$. The total entropy production is the sum of the $N$ yields:
@@
The fastest approach to reversible transport is achieved on geodesics. In the classical case, inserting the geodesic length (5) in (16), the ultimate lower bound on entropy production reads
$\Delta S=\frac{2}{N}(\arccos F)^{2}.$ (17)
This expresses the desired sharp operational meaning of the classical Bhattacharyya fidelity (1). The bounds are attainable in the limit $N\rightarrow\infty$ by operations on the system plus the reservoirs.
We find the simple meaning of the Fisher–KMB length if we introduce the number $\nu$ of equilibrations per unit length, i.e., the density $\nu=N/\ell_{\gamma}$. Then the entropy production becomes the following linear function of the path length:
@@
This suggests the following operational meaning of the Fisher–KMB length. Consider the near-reversible state transport via sequential equilibrations along a smooth path. If $\nu\gg 1$ is the density of equilibrations per unit path length then the entropy production is $(1/2\nu)$ per unit length.
@@ III Remarks and summary
Paths in Fisher geometries are perhaps the simplest non-local objects and, as such, they can be interpreted as state transport processes. It has therefore seemed logical to look here for the meaning of the geometries. One should mention the standard Kantorovich–Wasserstein theory Kantorovich (1960); Vaserstein (1969) where the transport from distribution $p$ to $q$ happens via direct relocation of populations from the $p_{\alpha}$’s into the $q_{\beta}$’s. The transport of quantum state $\hat{\rho}$ into $\hat{\sigma}$, too, can work similarly after a single unitary rotation of $\hat{\rho}$ that makes it commute with $\hat{\sigma}$. This may seem much simpler than sequential equilibration. But the simple method assumes the detailed control of the states in question. If it is not possible because, e.g., the states are complicated many-body states then the method of equilibration becomes more valuable.
@@
Thermodynamic systems are the typical examples. The Weinhold–Ruppeiner geometry Weinhold (1975); Ruppeiner (1979) defines the thermodynamic distances between thermodynamic states and the thermodynamic path lengths in general. It has long been known that the thermodynamic lengths coincide with statistical lengths Diósi et al. (1984) and they are rooted in the the Fisher geometry of Gibbs states Janyszek and Mrugała (1989). In thermodynamics, the expression $(\Delta\ell)^{2}/V$ quantifies how the statistical fluctuations of the thermodynamic parameters tend to zero in the thermodynamic limit of infinite volume $V\rightarrow\infty$. This does not interpret the global thermodynamic lengths $\ell$ but the local metrics. Studies of thermodynamic state transport Salamon and Nulton (1998); Salamon et al. (2023); Diósi (2002); Diosi et al. (1996) targeted a relationship between the thermodynamic path length and the thermodynamic entropy production. Salamon and Nulton Salamon and Nulton (1998) recognized that in discrete sequential equilibration the equilibration rates cancel from the expression of entropy production and they found the thermodynamic predecessor of eq. (16). Scandi and Perernau-Llobet derived similar result in quantum thermodynamics, using Fisher–KMB geometry of quantum Gibbs states Scandi and Perarnau-Llobet (2019). Ref. Diósi (2002) was the first attempt at constructing an underlying classical microscopic mechanism of a specific irreversible state transport. The present work constructed the abstract microscopic generalization valid in the Fisher–KMB geometries of arbitrary quantum states including the Fisher geometry special case of probability distributions. For this construction, the analytically tractable microscopic reservoir model and the related theorem, both proposed in ref. Diósi et al. (2006), have been instrumental.
@@
The global Fisher statistical distance both on space of probability distributions and on the space of density matrices with the KMB metrics got an operational interpretation for the first time. We obtain the sharp, not just qualitative, operational (physical) meaning of the classical Bhattacharyya fidelity. (Interestingly, the global Fisher–Bures geometry and the quantum fidelity are still waiting for their operational interpretations.) The concrete choice of the reservoir and the equilibration protocol is not likely to be critical and alternative microscopic models should lead to the same relationships between entropy production and the Fisher–KMB geometry. Future investigations may replace our stepwise transport with some continuous protocol.
@@ Acknowledgements.
The author thanks Florio M. Ciaglia who drew his attention to KMB metrics. This research was funded by the National Research, Development and Innovation Office (Hungary) Frontline Research Excellence Program (Grant No. KKP133827).